Exercise 2: Monopoly Assume a monopoly market with a demand curve of Q(p) = 10 – p and marginal cost, which are constant at 4Hi, Can anyone solve this question ?
Answer #1

& liven. 10-F Up) Mc = 4 A = 10 (all The is MR = MC The profit maximizing condition_undul mimoholy Marginal 1. Ma Revenu MargPage No. 4 MC D MR 5 * The 3 Quantity Ikue prefit maximizing facice =] pecefit._ maximizing quantity (6) Price Plasticity ofO Quantity under Regulated price Page Ne PR = brice undur brice Regulated > MC-4 . 10-4 Q 10- -) p = 10-B at of himnuun, 10-Q